OpenStudy (anonymous):

if p>=q>=5 where p and q are primes prove that 24|(p-q)(p+q)

OpenStudy (kinggeorge):

Since the primes are greater than 5, try dividing them up into primes congruent to 1 and 3 mod 4.

OpenStudy (kinggeorge):

Also into 1 and 5 mod 6.

OpenStudy (kinggeorge):

In the off-chance that p=q, then \(p-q=0\) and every number divides 0, so it still works.

myininaya (myininaya):

This is as far as I can get: \[24|(p-q)(p+q) => 24 \cdot k=(p-q)(p+q) \text{ for some integer k}\] p and q can't be even since p and q>=5 so p and q are odd primes p=2n+1 , q=2i+1, for some integers n>=2, i>=2 Now this gives us p-q=2(n-i) and p+q=2(n+i+1) So we have 24k=4(n-i)(n+i+1) => we really need to look at 6k=(n-i)(n+i+1)

OpenStudy (kinggeorge):

Look at it this way. Suppose \(p\equiv q \equiv 1 \mod 4\) and \(p\equiv q \equiv 1 \mod 6\). Then \(12|(p-q)\) (proof left to the reader). However, \(2|(p+q)\) so 24 divides their product.

OpenStudy (kinggeorge):

Also, if \(p \equiv 1 \mod 4\) and \(q \equiv 3 \mod 4\) and \(p \equiv q\equiv 1 \mod 6\). Then \(4|(p+q)\) and \(6|(p-q)\). Thus \(24|(p-q)(p+q)\).

OpenStudy (kinggeorge):

By continuing these different cases you'll find that it's always true.